# SoRA: Singular Value Decomposed Low-Rank Adaptation for Domain Generalizable Representation Learning

### v1

@@ Abstract
11footnotetext: Corresponding author.Domain generalization (DG) aims to adapt a model using one or multiple source domains to ensure robust performance in unseen target domains. Recently, Parameter-Efficient Fine-Tuning (PEFT) of foundation models has shown promising results in the context of DG problem. Nevertheless, existing PEFT methods still struggle to strike a balance between preserving generalizable components of the pre-trained model and learning task-specific features. To gain insights into the distribution of generalizable components, we begin by analyzing the pre-trained weights through the lens of singular value decomposition. Building on these insights, we introduce Singular Value Decomposed Low-Rank Adaptation (SoRA), an approach that selectively tunes minor singular components while keeping the residual parts frozen. SoRA effectively retains the generalization ability of the pre-trained model while efficiently acquiring task-specific skills. Furthermore, we freeze domain-generalizable blocks and employ an annealing weight decay strategy, thereby achieving an optimal balance in the delicate trade-off between generalizability and discriminability. SoRA attains state-of-the-art results on multiple benchmarks that span both domain generalized semantic segmentation to domain generalized object detection. In addition, our methods introduce no additional inference overhead or regularization loss, maintain compatibility with any backbone or head, and are designed to be versatile, allowing easy integration into a wide range of tasks.
@@ 1 Introduction
To uncover the structure in which world knowledge is encoded within the pre-trained weights, we first perform singular value decomposition (SVD) on the learned weight matrices and then analyze the behavior of VFMs after selectively removing specific singular components111In this paper, we refer to the combination of a singular value and its corresponding singular vectors as a singular component.. Our analysis reveals that singular vectors associated with higher singular values tend to extract general features spanning multiple classes in ImageNet, while minor singular components are primarily responsible for capturing context-specific features. Capitalizing on our findings, we introduce Singular Value Decomposed Low-Rank Adaptation (SoRA), which starts by performing SVD to decompose the weights into their singular vectors, i.e., , then fine-tunes only the minor singular components. Specifically, we initialize LoRA adapter using the components with the smallest singular values, i.e., , and the remaining residual components are frozen to maintain generalizability, as illustrated in Fig. 1 right.
By extending our analysis from the weight-level to the block-level, we observe and empirically demonstrate that the early blocks of VFMs effectively extract well-localized semantic features and are less affected by style or input-level domain shifts (see Fig. 3). Therefore, we freeze these blocks to preserve generalizable components while reducing the number of trainable parameters. Although our proposed methods are effective, tuning the less-optimized bottom spectral space and freezing early blocks may lead to a lack of discriminability. To mitigate this issue, we introduce an annealing weight decay scheme that gradually reduces the regularization loss incurred by weight decay over the course of training. As shown in Fig. 1, comprehensive evaluations on DG benchmarks demonstrate that our framework outperforms previous state-of-the-art baselines and FFT, training only 0.58% to 1.6% of the parameters. Furthermore, SoRA demonstrates impressive scaling performance with both data and model size.
@@
We present SoRA, a novel PEFT method specifically tailored for DG problem, which leverages spectral information to retain the generalization capacity of VFMs. In addition, we identify and freeze domain-generalizable blocks while employing an annealing weight decay strategy, thereby achieving an optimal balance between generalizability and discriminability.
Extensive experiments on various DG for semantic segmentation and object detection tasks reveal that our framework surpasses existing baselines by a substantial margin. Notably, SoRA achieves an mIoU of 80.4% on Cityscapes without accessing real-world datasets, while training only 0.58% of the parameters. Additionally, under the most challenging daytime-sunny night-rainy detection setting, SoRA outperforms the prior SOTA implementation by achieving a +23.5% mAP gain.
@@ 2 Related Work
Vision foundation models (VFMs). Recently, the advent of VFMs with general perception capabilities has laid the foundation for a paradigm in which these models are broadly applicable to numerous downstream tasks. Seminal examples of models advancing this direction include CLIP [65], which demonstrates strong zero-shot generalizability through training on web-scale, weakly supervised image-text pairs; EVA02 [24], which enhances CLIP features with masked image modeling [31]; and DINOv2 [58], which incorporates losses from prior arts like iBoT [92] and DINO [8], and is trained on massive curated datasets, offering strong spatial features for dense prediction tasks. Our SoRA builds on this foundation, harnessing the full potential of VFMs in the context of DG problem.
@@ 4 Proposed Method
4.1 Singular Value Decomposed Low-Rank Adaptation
To effectively preserve the integrity of generalizable representations in pre-trained weights, we propose Singular value decomposed Low-Rank Adaptation (SoRA). Firstly, SoRA performs singular value decomposition (SVD) on the pre-trained weight matrices, including those from the attention layers, multi-layer perceptron (MLP) layers, and more. The weight matrix is decomposed using SVD as follows:
@@ 4.1 Singular Value Decomposed Low-Rank Adaptation
Importantly, SoRA offers several key advantages for tackling the DG problem: i) SoRA initializes the adapter using components orthogonal to the principal components, thus minimizing interference with pre-trained representations. For the batched version of the forward pass described in Eq. 1 (with and ), the gradients of and are and . Therefore, unlike LoRA, which randomly initializes , our approach facilitates convergence along singular directions that preserve the integrity of the generalizable components. ii) Consistent with prior DG methods [14, 45, 43, 91] that guide models to prioritize shape over texture, thereby achieving superior out-of-distribution performance, SoRA effectively preserves the domain generalizable shape-biased features of VFMs by tuning only the minor singular components (see Fig. 2 bottom and Tab. 1).
@@
For ease of analysis, we partition the 1024 singular directions into four equal groups, calculate the average SMR for each group, and then compare the resulting values between SoRA and LoRA. As shown in Tab. 1, the SoRA initialization scheme proves more effective than LoRA in minimizing interference with singular components with higher singular values. In summary, our method effectively preserves the structure of the VFM’s generalizable representations while adapting to downstream tasks, resulting in significant improvements on DG benchmarks compared to SOTA baselines, as confirmed by our experimental results.
@@ 4.3 Annealing Weight Decay
Our methods are specialized in preserving the pre-trained knowledge of VFMs, but this may come at the cost of diminished discriminability. One can remove weight decay or set its coefficient to a very small value to enhance discriminability. However, doing so risks increased interference of the tuned parameters with the principal components of the pre-trained weights. Therefore, based on the observation that regularization primarily affects the early phases of learning by guiding a model towards having good generalization properties [26], we leverage an Annealing Weight Decay (AWD) scheme. Specifically, AWD starts with a relatively large weight decay coefficient and progressively reduces it to zero as training progresses. By default, we use a cosine schedule for AWD. As shown in Tab. 1 ( vs. ) and Tab. 10, AWD integrates seamlessly with SoRA, improving discriminability without compromising SoRA’s ability to preserve pre-trained knowledge.
@@ 5 Experiments
We conduct a variety of experiments to showcase the efficacy of SoRA on DG benchmarks including domain generalized semantic segmentation (DGSS) and domain generalized object detection (DGOD). We then perform an ablation study on SoRA, emphasizing the contribution of each component to the DG performance. Lastly, we extend our approach beyond dense predictions to generative modeling.
@@ 5.1 Experimental Setup
| SoRA (Ours) | EVA02-L | 5.1M | 68.05 | 60.81 | 68.33 | 65.73 |
| SoRA (Ours) | EVA02-L | 5.1M | 69.94 | 62.48 | 68.33 | 66.92 |
@@
| SoRA (Ours) | DINOv2-L | 4.9M | 71.82 | 61.31 | 71.67 | 68.27 |
| SoRA (Ours) | DINOv2-L | 4.9M | 73.63 | 63.33 | 70.98 | 69.31 |
@@
| SoRA (Ours) | DINOv2-L | 4.9M | 73.16 | 61.90 | 72.73 | 69.26 |
| SoRA (Ours) | DINOv2-L | 4.9M | 74.85 | 63.59 | 73.92 | 70.79 |
@@
| SoRA (Ours) | DINOv2-L | 4.9M | 77.33 | 62.78 | 74.93 | 71.68 |
@@
| SoRA (Ours) | DINOv2-G | 6.6M | 78.39 | 63.75 | 75.16 | 72.43 |
| SoRA (Ours) | DINOv2-G | 6.6M | 80.37 | 65.67 | 76.18 | 74.07 |
@@
Implementation details. Since our method focuses on preserving pre-trained knowledge, we mainly evaluate its effectiveness using VFMs like EVA02 [24] and DINOv2 [58] as the backbone. To further elevate performance, we leverage state-of-the-art decode heads such as Mask2Former [13] and Co-DETR [93]. Extensive experiments with various backbones and heads are provided in the Supplemental. By default, SoRA is applied to all linear layers within the self-attention and MLP, with a rank of 16. Detailed training settings are presented in the Supplemental.
@@ 5.2 Comparison with State-of-the-Arts
We compare SoRA with previous SOTA baselines, achieving the highest performance across all DG scenarios without adding inference latency or regularization loss. In our tables, indicates trainable parameters in the backbones. Marker refers to models tested with multi-scale flip augmentation. represents re-implemented test results using official checkpoints for a fair comparison, and denotes training on images cropped to 10241024; otherwise, the default is 512512. A comprehensive comparison with earlier works can be found in the Supplemental.
@@
| SoRA (Ours) | DINOv2-L | 4.9M | 67.02 | 76.45 | 71.74 |
| SoRA (Ours) | DINOv2-L | 4.9M | 68.08 | 77.87 | 72.98 |
@@
Synthetic-to-real DGSS. In Tab. 2, we adopt GTAV {Citys., BDD, Map.} as the basic setting, and demonstrate scalability of SoRA by incorporating additional synthetic datasets into the training process. We report the mean Intersection over Union (mIoU) and benchmark our results against SOTA methods. SoRA consistently outperforms VFM-based DGSS methods and PEFT approaches using both vision-language (EVA02) and vision-only (DINOv2) VFM backbones. Notably, SoRA achieves superior results over LoRA by minimally interfering with the hierarchical structure of VFM’s pre-trained knowledge (see Tab. 1). Moreover, SoRA exhibits superior scalability across data, model, and input size, with improvements becoming more pronounced when test-time augmentation is applied.
Real-to-real DGSS. In Tables 5 and 5, we conduct experiments under the Citys. {BDD, Map.} / ACDC settings. When combined with DINOv2-Large backbone, SoRA consistently gains the best results across all settings. This highlights the effectiveness of our method as a strong solution for real-world deployment scenarios, where domain shifts caused by geographic or weather variations are common.
@@
| SoRA (Ours) | 4.9M | 69.4 | 59.3 | 59.3 | 47.6 | 51.0 | 54.3 |
| Efficiency | Training (bs = 4) | Inference (bs = 1 / 32) | ||
| Methods | Time (hrs) | Memory | Throughput (imgs/s) | Memory |
| 9.2 | 12.5G | 20.0 / - | 5.5G / OOM | |
| 9.3 | 12.2G | 33.6 / 64.3 | 4.7G / 48.2G | |
| 9.0 | 12.7G | 56.4 / 79.7 | 4.4G / 40.9G | |
| 27.9 | 45.3G | 21.6 / - | 10.6G / OOM | |
| 18.9 | 25.6G | 21.6 / - | 10.6G / OOM | |
Real-world data efficiency. Following [80], we evaluate models under constrained conditions using a limited set of real-world images. Specifically, models pre-trained on synthetic datasets GTAV + SYNTHIA + UrbanSyn (marked with in Tab. 2) are fine-tuned on 1/16 of the Cityscapes training set. As shown in Tab. 5, SoRA achieves performance on par with Rein in the source domain, while offering greater robustness on unseen domains. This underscores SoRA’s data-efficient adaptability to real-world scenes, while effectively avoiding overfitting to the source domain.
Clear-to-adverse weather DGOD. In Tab. 7, we compare SoRA with existing DGOD methods for classic backbone-head setting and with PEFT methods for recent foundation architectures. As expected, SoRA demonstrates only minor improvements in the classic setting; however, in the VFM setting—where there is considerably more knowledge to retain—it shows substantial performance gains. As illustrated by the visualization results in Fig. 4, SoRA makes robust predictions even under diverse challenging scenes.
@@
Model efficiency. As detailed in Tab. 7, SoRA exhibits higher throughput than adapter- and VPT-based methods like Rein [80] and SET [86], as it incurs no additional latency. This advantage is especially significant in online inference settings, where the batch size is typically as small as one [33]. Furthermore, in scenarios involving DINOv2-giant exceeding 1B parameters, SoRA can drastically reduce training costs compared to FFT. SoRA initialization is completed within 30 seconds for large-sized models, which is a negligible cost given the improved performance.
@@ 5.3 Ablation Study
Component Analysis. A thorough examination of each component of SoRA in the DGOD and DGSS basic setting, as summarized in Tab. 10, demonstrates that all components of SoRA incrementally improve DG performance without adding computational costs. Notably, the marked performance gap with tuning principal singular components [55] suggests that SoRA better preserves the integrity of pre-trained representations while learning task-specific features.
Tuning granularity. Interestingly, as shown in Tables 10 and 10, our method consistently achieves competitive results compared to SOTA baselines across all tuning granularities, indicating that SoRA can be flexibly configured according to the training budget, task difficulty, or domain gap. The best results are achieved with the default setting.

### v2

# SoMA: Singular Value Decomposed Minor Components Adaptation for Domain Generalizable Representation Learning

@@ Abstract
11footnotetext: Corresponding author.Domain generalization (DG) aims to adapt a model using one or multiple source domains to ensure robust performance in unseen target domains. Recently, Parameter-Efficient Fine-Tuning (PEFT) of foundation models has shown promising results in the context of DG problem. Nevertheless, existing PEFT methods still struggle to strike a balance between preserving generalizable components of the pre-trained model and learning task-specific features. To gain insights into the distribution of generalizable components, we begin by analyzing the pre-trained weights through the lens of singular value decomposition. Building on these insights, we introduce Singular Value Decomposed Minor Components Adaptation (SoMA), an approach that selectively tunes minor singular components while keeping the residual parts frozen. SoMA effectively retains the generalization ability of the pre-trained model while efficiently acquiring task-specific skills. Moreover, we freeze domain-generalizable blocks and employ an annealing weight decay strategy, thereby achieving an optimal balance in the delicate trade-off between generalizability and discriminability. SoMA attains state-of-the-art results on multiple benchmarks that span both domain generalized semantic segmentation to domain generalized object detection. In addition, our methods introduce no additional inference overhead or regularization loss, maintain compatibility with any backbone or head, and are designed to be versatile, allowing easy integration into a wide range of tasks.
@@ 1 Introduction
To uncover the structure in which world knowledge is encoded within the pre-trained weights, we first perform singular value decomposition (SVD) on the learned weight matrices and then analyze the behavior of VFMs after selectively removing specific singular components111In this paper, we refer to the combination of a singular value and its corresponding singular vectors as a singular component.. Our analysis reveals that singular vectors associated with higher singular values tend to extract general features spanning multiple classes in ImageNet, while minor singular components are primarily responsible for capturing context-specific features. Capitalizing on our findings, we introduce Singular Value Decomposed Minor Components Adaptation (SoMA), which starts by performing SVD to decompose the weights into their singular vectors, i.e., , then fine-tunes only the minor singular components. Specifically, we initialize LoRA adapter using the components with the smallest singular values, i.e., , and the remaining residual components are frozen to maintain generalizability, as illustrated in Fig. 1 right.
By extending our analysis from the weight-level to the block-level, we observe and empirically demonstrate that the early blocks of VFMs effectively extract well-localized semantic features and are less affected by style or input-level domain shifts (see Fig. 3). Therefore, we freeze these blocks to preserve generalizable components while reducing the number of trainable parameters. Although our proposed methods are effective, tuning the less-optimized bottom spectral space and freezing early blocks may lead to a lack of discriminability. To mitigate this issue, we introduce an annealing weight decay scheme that gradually reduces the regularization loss incurred by weight decay over the course of training. As shown in Fig. 1, comprehensive evaluations on DG benchmarks demonstrate that our framework outperforms previous state-of-the-art baselines and FFT, training only 0.58% to 1.6% of the parameters. Furthermore, SoMA demonstrates impressive scaling performance with both data and model size.
@@
We present SoMA, a novel PEFT method specifically tailored for DG problem, which leverages spectral information to retain the generalization capacity of VFMs. In addition, we identify and freeze domain-generalizable blocks while employing an annealing weight decay strategy, thereby achieving an optimal balance between generalizability and discriminability.
Extensive experiments on various DG for semantic segmentation and object detection tasks reveal that our framework surpasses existing baselines by a substantial margin. Notably, SoMA achieves an mIoU of 80.4% on Cityscapes without accessing real-world datasets, while training only 0.58% of the parameters. Additionally, under the most challenging daytime-sunny night-rainy detection setting, SoMA outperforms the prior SOTA implementation by achieving a +23.5% mAP gain.
@@ 2 Related Work
Vision foundation models (VFMs). Recently, the advent of VFMs with general perception capabilities has laid the foundation for a paradigm in which these models are broadly applicable to numerous downstream tasks. Seminal examples of models advancing this direction include CLIP [65], which demonstrates strong zero-shot generalizability through training on web-scale, weakly supervised image-text pairs; EVA02 [24], which enhances CLIP features with masked image modeling [31]; and DINOv2 [58], which incorporates losses from prior arts like iBoT [92] and DINO [8], and is trained on massive curated datasets, offering strong spatial features for dense prediction tasks. Our SoMA builds on this foundation, harnessing the full potential of VFMs in the context of DG problem.
@@ 3 Preliminaries
Problem Formulation. The entire model consists of a feature extractor , which combines a VFM parameterized by with our SoMA adapter parameterized by , and a task-specific decode head , parameterized by . The optimization objective can be expressed as follows:
where () denotes an input image and its associated ground truth, drawn from the source domain datasets , comprising samples. Our proposed SoMA framework facilitates generalizing task-specific skills learned from to the unseen target domain datasets .
@@ 4 Proposed Method
4.1 Minor Singular Components Adaptation
To effectively preserve the integrity of generalizable representations in pre-trained weights, we propose Singular value decomposed Minor components Adaptation (SoMA). Firstly, SoMA performs singular value decomposition (SVD) on the pre-trained weight matrices, including those from the attention layers, multi-layer perceptron (MLP) layers, and more. The weight matrix is decomposed using SVD as follows:
@@ 4.1 Minor Singular Components Adaptation
Importantly, SoMA offers several key advantages for tackling the DG problem: i) SoMA initializes the adapter using components orthogonal to the principal components, thus minimizing interference with pre-trained representations. For the batched version of the forward pass described in Eq. 1 (with and ), the gradients of and are and . Therefore, unlike LoRA, which randomly initializes , our approach facilitates convergence along singular directions that preserve the integrity of the generalizable components. ii) Consistent with prior DG methods [14, 45, 43, 91] that guide models to prioritize shape over texture, thereby achieving superior out-of-distribution performance, SoMA effectively preserves the domain generalizable shape-biased features of VFMs by tuning only the minor singular components (see Fig. 2 bottom and Tab. 1).
@@
For ease of analysis, we partition the 1024 singular directions into four equal groups, calculate the average SMR for each group, and then compare the resulting values between SoMA and LoRA. As shown in Tab. 1, the SoMA initialization scheme proves more effective than LoRA in minimizing interference with singular components with higher singular values. In summary, our method effectively preserves the structure of the VFM’s generalizable representations while adapting to downstream tasks, resulting in significant improvements on DG benchmarks compared to SOTA baselines, as confirmed by our experimental results.
@@ 4.3 Annealing Weight Decay
Our methods are specialized in preserving the pre-trained knowledge of VFMs, but this may come at the cost of diminished discriminability. One can remove weight decay or set its coefficient to a very small value to enhance discriminability. However, doing so risks increased interference of the tuned parameters with the principal components of the pre-trained weights. Therefore, based on the observation that regularization primarily affects the early phases of learning by guiding a model towards having good generalization properties [26], we leverage an Annealing Weight Decay (AWD) scheme. Specifically, AWD starts with a relatively large weight decay coefficient and progressively reduces it to zero as training progresses. By default, we use a cosine schedule for AWD. As shown in Tab. 1 ( vs. ) and Tab. 9, AWD integrates seamlessly with SoMA, improving discriminability without compromising SoMA’s ability to preserve pre-trained knowledge.
@@ 5 Experiments
We conduct a variety of experiments to showcase the efficacy of SoMA on DG benchmarks including domain generalized semantic segmentation (DGSS) and domain generalized object detection (DGOD). We then perform an ablation study on SoMA, emphasizing the contribution of each component to the DG performance. Lastly, we extend our approach beyond dense predictions to generative modeling.
@@ 5.1 Experimental Setup
| SoMA (Ours) | EVA02-L | 5.1M | 68.05 | 60.81 | 68.33 | 65.73 |
| SoMA (Ours) | EVA02-L | 5.1M | 69.94 | 62.48 | 68.33 | 66.92 |
@@
| SoMA (Ours) | DINOv2-L | 4.9M | 71.82 | 61.31 | 71.67 | 68.27 |
| SoMA (Ours) | DINOv2-L | 4.9M | 73.63 | 63.33 | 70.98 | 69.31 |
@@
| SoMA (Ours) | DINOv2-L | 4.9M | 73.16 | 61.90 | 72.73 | 69.26 |
| SoMA (Ours) | DINOv2-L | 4.9M | 74.85 | 63.59 | 73.92 | 70.79 |
@@
| SoMA (Ours) | DINOv2-L | 4.9M | 77.33 | 62.78 | 74.93 | 71.68 |
@@
| SoMA (Ours) | DINOv2-G | 6.6M | 78.39 | 63.75 | 75.16 | 72.43 |
| SoMA (Ours) | DINOv2-G | 6.6M | 80.37 | 65.67 | 76.18 | 74.07 |
@@
Implementation details. Since our method focuses on preserving pre-trained knowledge, we mainly evaluate its effectiveness using VFMs like EVA02 [24] and DINOv2 [58] as the backbone. To further elevate performance, we leverage state-of-the-art decode heads such as Mask2Former [13] and Co-DETR [93]. Extensive experiments with various backbones and heads are provided in the Supplemental. By default, SoMA is applied to all linear layers within the self-attention and MLP, with a rank of 16. Detailed training settings are presented in the Supplemental.
@@ 5.2 Comparison with State-of-the-Arts
We compare SoMA with previous SOTA baselines, achieving the highest performance across all DG scenarios without adding inference latency or regularization loss. In our tables, indicates trainable parameters in the backbones. Marker refers to models tested with multi-scale flip augmentation. represents re-implemented test results using official checkpoints for a fair comparison, and denotes training on images cropped to 10241024; otherwise, the default is 512512. A comprehensive comparison with earlier works can be found in the Supplemental.
@@
| SoMA (Ours) | DINOv2-L | 4.9M | 67.02 | 76.45 | 71.74 |
| SoMA (Ours) | DINOv2-L | 4.9M | 68.08 | 77.87 | 72.98 |
@@
Synthetic-to-real DGSS. In Tab. 2, we adopt GTAV {Citys., BDD, Map.} as the basic setting, and demonstrate scalability of SoMA by incorporating additional synthetic datasets into the training process. We report the mean Intersection over Union (mIoU) and benchmark our results against SOTA methods. SoMA consistently outperforms VFM-based DGSS methods and PEFT approaches using both vision-language (EVA02) and vision-only (DINOv2) VFM backbones. Notably, SoMA achieves superior results over LoRA by minimally interfering with the hierarchical structure of VFM’s pre-trained knowledge. Moreover, SoMA exhibits superior scalability across data, model, and input size, with improvements becoming more pronounced when test-time augmentation is applied. We provide comparisons regarding model efficiency in the Supplemental.
Real-to-real DGSS. In Tables 5 and 5, we conduct experiments under the Citys. {BDD, Map.} / ACDC settings. When combined with DINOv2-L backbone, SoMA consistently gains the best results across all settings. This highlights the effectiveness of our method as a strong solution for real-world deployment scenarios, where domain shifts caused by geographic or weather variations are common.
@@
| SoMA (Ours) | 4.9M | 69.4 | 59.3 | 59.3 | 47.6 | 51.0 | 54.3 |
Real-world data efficiency. Following [80], we evaluate models under constrained conditions using a limited set of real-world images. Specifically, models pre-trained on synthetic datasets GTAV + SYNTHIA + UrbanSyn (marked with in Tab. 2) are fine-tuned on 1/16 of the Cityscapes training set. As shown in Tab. 5, SoMA achieves performance on par with Rein in the source domain, while offering greater robustness on unseen domains. This underscores SoMA’s data-efficient adaptability to real-world scenes, while effectively avoiding overfitting to the source domain.
Clear-to-adverse weather DGOD. In Tab. 6, we compare SoMA with existing DGOD methods for classic backbone-head setting and with PEFT methods for recent foundation architectures. As expected, SoMA demonstrates only minor improvements in the classic setting; however, in the VFM setting—where there is considerably more knowledge to retain—it shows substantial performance gains. As illustrated by the visualization results in Fig. 4, SoMA makes robust predictions even under diverse challenging scenes.
@@ 5.3 Ablation Study
Component Analysis. A thorough examination of each component of SoMA in the DGOD and DGSS basic setting, as summarized in Tab. 9, demonstrates that all components of SoMA incrementally improve DG performance without adding computational costs. Notably, the marked performance gap with tuning principal singular components [55] suggests that SoMA better preserves the integrity of pre-trained representations while learning task-specific features.
Tuning granularity. Interestingly, as shown in Tables 9 and 9, our method consistently achieves competitive results compared to SOTA baselines across all tuning granularities, indicating that SoMA can be flexibly configured according to the training budget, task difficulty, or domain gap. The best results are achieved with the default setting.